Joel spent 15hrs watching movies. The space adventures were 1.5hrs long and the mysteries were 2hrs long. He watched twice as many space adventures as mysteries. How much time did he spend watching space adventures?
Found 2 solutions by amoresroy, mananth:
Answer by amoresroy(361)

Joel spent 15hrs watching movies. The space adventures were 1.5hrs long and the mysteries were 2hrs long. He watched twice as many space adventures as mysteries. How much time did he spend watching space adventures?
Let 1.5s = time spent on watching space adventures
2m = time spent on watching mysteries
Equation 1
1.5s + 2m = 15
Equation 2
2m = s
Substitute 2m = s in equation 1
1.5s + s = 15
2.5s = 15
s = 6
1.5s = 1.5(6) = 9
Joel spent 9 hours watching space adventures




Answer by mananth(16949)

let mysteries movies watched be x
space adventures movies watched will be 2x
..
mysteries = 2hours long
space adventures = 1.5 hours
..
2x+1.5*2x =15
2x+3x=15
5x=15
/5
x= 3 movies mysteries
6 movies space adventure
6*1.5 = 9 hours of space adventure
...
m.ananth@hotmail.ca